Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2018-10-04

Observers: T.Reynolds, S.Moran
Support Team: J. Harmanen, H. Kuncarayakti
TAT: Joe Anderson, Lluís Galbany

Calibrations: 
Dome flats and arcs for gr11, 13, 16, 20, biases

Weather: 
Start - Very cloudy (closed)
Middle - Clouds
End - Clouds

UT       Target       Type      Setup          Exp         airmass  DIMM  note
-----------------------------------------------------------------------------------------
                                             
EFOSC                                                                       
Stayed closed at the beginning of the night due to thick clouds
Opened at 02:30
02:30    Image analysis
02:50    2018gft    follow-up  gr16/1.5"+1.0",gr11/1.0" 2400 1-1.10 ~1.1
05:45    SOFI service mode observations
06:45    Image analysis
07:00    2016ezh    follow-up  gr20,gr13/1.0" 1800*2,1500  1.2-1.70 ~1.4
09:00    gd71       std        gr11/13/16/20/1.0"  100     1.4      ~1.3


Comments